# Universidad de Alcalá, Alcalá de Henares 28805, Spain

### v1

@@ Abstract
Aims. We aim to investigate the origin and transport of energetic particles in two rare bidirectional anisotropic SEP events observed by Solar Orbiter, with particular emphasis on their association with magnetic flux ropes.
Methods. SEP, solar wind plasma, magnetic field, and solar radio measurements were analyzed. Velocity dispersion analysis determined release times and path lengths for distinct particle populations. Automated flux rope identification and magnetic helicity diagnostics were used to characterize magnetic flux ropes.
Results. Both events showed two clear velocity dispersion signatures with opposite particle anisotropies during their onset phase. The sunward streaming protons, characterized by delayed release time, harder spectral index, and higher intensities, may be attributed to coronal mass ejection-driven shock acceleration, while the promptly released anti-sunward streaming protons are likely linked to flare acceleration. Notably, in both cases, small-scale flux ropes were identified in situ during the time intervals corresponding to the bidirectional particle streaming. Path lengths derived for sunward and anti-sunward injections were substantially greater than nominal values of the Parker field lines, further supporting the role of the flux rope in shaping particle trajectories.
Conclusions. These observations demonstrate that magnetic flux rope could significantly affect magnetic connectivity to the source region and SEP propagation in the inner heliosphere, while simultaneous velocity dispersion from two distinct particle sources allows for direct constraints on the topology of the flux rope. Our results highlight the value of combining particle anisotropy, release time, source spectra, and magnetic structure diagnostics to unravel SEP transport in complex transient magnetic structures, and also present new challenges for the current SEP transport model.
@@
Solar Energetic Particle (SEP) events are characterized by transient enhancements in the intensity of energetic charged particles, which are accelerated from the solar corona during eruptive phenomena such as flares and coronal mass ejections (CMEs) and subsequently propagate along magnetic field lines into interplanetary space (e.g., Desai & Giacalone, 2016). The particle anisotropy measured at the spacecraft serves as a direct indicator of particle transport processes. Specifically, large anisotropy is typically associated with weak pitch angle scattering, meaning that particle flux is strongly peaked along certain magnetic field directions, with most particles streaming either sunward or anti-sunward. In contrast, small or near-zero anisotropy suggests strong pitch angle scattering or enhanced perpendicular diffusion, resulting in a more isotropic particle distribution (Dresing et al., 2014). Therefore, anisotropy is a crucial diagnostic for identifying atypical SEP transport processes, especially in complex interplanetary environments.
@@
While most SEP events exhibit outward streaming particles along open magnetic field lines, an increasing number of in-situ observations have revealed rare cases of sunward-directed SEP events, where particles are detected propagating back toward the Sun with high intensity (e.g., Richardson & Cane, 1996; Larson et al., 1997; Rodriguez-Pacheco, 1999; Richardson et al., 2000; Malandraki et al., 2002; Rodríguez-Pacheco et al., 2003; Leske et al., 2012; Tan et al., 2012; Gómez-Herrero et al., 2017; Li et al., 2020; Wei et al., 2024; Rodríguez-García et al., 2025). These events typically involve special interplanetary conditions, such as an observer being embedded within a magnetic cloud (MC), the presence of a shock or compressed magnetic fields acting as a magnetic mirror, or remote acceleration sites away from the observer. Based on previous case studies, sunward SEP events can be generally categorized into four main types according to their underlying physical mechanisms. Firstly, the observer may reside within a closed magnetic flux rope (e.g., a magnetic cloud) with both legs anchored at the Sun, enabling bidirectional (sunward and anti-sunward) particle flows via dual injections from two legs or reflection at the footpoint (Dresing et al., 2016; Gómez-Herrero et al., 2017; Rodríguez-García et al., 2025). Moreover, the three-dimensional topology of the flux rope also allows for the possibility that particles are injected from the north or south directions (Wimmer-Schweingruber et al., 2023). Secondly, sunward SEPs can arise from magnetic mirroring, where SEPs transport along open field lines and encounter a region of enhanced magnetic field strength, such as a shock or a CME sheath, which acts as a magnetic mirror and reflects particles toward the Sun (Malandraki et al., 2002; Klassen et al., 2012; Tan et al., 2012; Li et al., 2020). In contrast to cases where an observer inside a closed flux rope may detect SEPs from different sources at each footpoint, the mirror effect results in the observation of particles originating from the same source. Thirdly, Wei et al. (2024) proposed that a remote strong accelerator (e.g., an interplanetary CME-driven shock ) located beyond the observer may inject SEPs that then propagate sunward along open magnetic field lines to the spacecraft. Finally, cross-field diffusion and field-line meandering may allow sunward SEPs to reach observers with poor magnetic connectivity to the source region (e.g., He, 2015). However, this mechanism alone is insufficient to explain sunward-propagating events observed at locations with good magnetic connectivity to the source.
@@
While previous studies have established several categories of sunward-propagating SEP events, here we investigate bidirectional SEP events that exhibit two distinct velocity dispersion signatures during their onset phases, with one associated with anti-sunward propagation and the other with sunward propagation. Such events, although rare, can provide critical constraints on both the path lengths and release times of different particle sources, and thereby offer unique insights into the large-scale configuration of the interplanetary magnetic field, as well as the particle acceleration and transport. In this work, we present a detailed analysis of two bidirectional, anisotropic SEP events observed by Solar Orbiter, each displaying two distinct velocity dispersion signatures and both likely associated with small-scale flux ropes (SFRs). To our knowledge, such events have not been previously reported. SFRs typically have durations ranging from tens of minutes to several hours. The origin of SFRs remains a subject of debate, as they are believed to result from solar eruptive processes, possibly as manifestations of small coronal mass ejections (Feng et al., 2007), and from in-situ formation in the heliosphere through magnetic reconnection (Cartwright & Moldwin, 2010). Notably, SFRs exhibit much shorter durations and lack the typical plasma signatures of large-scale CMEs (e.g., Wimmer-Schweingruber et al., 2006; Hu et al., 2018). As a result, SFRs are more challenging to detect and interpret, yet they could play an important role in modulating energetic particle propagation in the solar wind.
@@
Through a combination of energetic particle measurements, velocity dispersion analysis (VDA), and magnetic field diagnostics, we investigate the origins, transport process and associated magnetic structure of two bidirectional SEP events. Section 2 describes the data and methods employed. The observations and interpretation of the two events are presented in Section 3. Finally, Section 4 summarizes the main findings and discusses their implications for understanding SEP transport and the topology of flux ropes.
@@ 2 Data and Methods
This study utilizes data collected by multiple instruments on board the Solar Orbiter. Specifically, SEP measurements are provided by the Energetic Particle Detector suite (EPD; Rodríguez-Pacheco et al. 2020; Wimmer-Schweingruber et al. 2021), which comprises the Supra-Thermal Electron and Proton sensor (STEP), the Electron Proton Telescope (EPT), the High Energy Telescope (HET), and the Suprathermal Ion Spectrograph (SIS). These instruments cover ion energies from a few keV nucleon-1 up to above 100 MeV nucleon-1. In this work, we mainly utilize ion measurements from EPT and HET, each offering four viewing directions, thereby enabling the analysis of particle anisotropy. While EPT does not explicitly differentiate between ion species, the measured fluxes are commonly interpreted as proton-dominated, since protons constitute the most abundant ion population in large SEP events. Magnetic field data are obtained from the Solar Orbiter Magnetometer (MAG; Horbury et al. 2020). Plasma parameters are measured with the Proton-Alpha Sensor (PAS), and the pitch angle distributions of suprathermal electrons are provided by the Electron Analyser System (EAS), both part of the Solar Wind Analyzer (SWA; Owen et al. 2020). Additionally, we examine radio data from the Radio and Plasma Waves (RPW; Maksimovic et al. 2020) instrument onboard Solar Orbiter.
@@
A key focus of this work is to investigate the roles of small-scale magnetic flux ropes in the bidirectional SEP events. Identifying SFRs in the solar wind presents a significant challenge due to their short durations, weaker coherent magnetic signatures compared to magnetic clouds, and ambiguous boundaries, especially in the presence of solar wind fluctuations. To address this, we employ the automated detection algorithm implemented in the PyGS package (Hu et al., 2018; Zheng & Hu, 2018; Chen & Hu, 2022), which leverages the unique characteristics of the Grad–Shafranov (GS) equation (Sonnerup & Guo, 1996; Hu & Sonnerup, 2002). In the local flux rope frame, a spacecraft crossing an SFR typically observes a bipolar rotation in one of the magnetic field components, corresponding to nested isosurfaces of the magnetic flux function . For a true flux rope, the transverse pressure () is expected to be a single-valued function of , resulting in a characteristic double-folding pattern of between the inbound and outbound traversals. The PyGS algorithm systematically scans in-situ data for intervals that exhibit this double-folding signature, enabling robust and automated identification of SFRs.
@@
However, in highly dynamic or turbulent environments, conventional automated methods such as PyGS may fail to reliably identify flux rope structures due to strong magnetic field fluctuations. To address these limitations, we supplement our analysis with calculations of the normalized reduced magnetic helicity, , which serves as a sensitive proxy for magnetic field twist. Following Zhao et al. (2020, 2021b), the scale- and time-dependent normalized reduced magnetic helicity, , is then evaluated as
@@
where the tilde represents wavelet-transformed quantities using the complex Morlet wavelet function (Torrence & Compo, 1998), refers to the imaginary part, represents the wavelet scale, and the asterisk denotes complex conjugation. Here, , , and correspond to the magnetic field components in the Radial-Tangential-Normal (RTN) coordinate system. As demonstrated in previous studies (Zhao et al., 2020, 2021a, 2021b), flux ropes are typically characterized by high values of magnetic helicity, which result from the large-angle rotation of the magnetic field. In this work, we use regions with as proxies for the possible presence of SFRs. Since strongly Alfvénic fluctuations also typically appear as intermittent enhancements of , to exclude their influence, we further examine the normalized cross helicity and the normalized residual energy , as described in Appendix A.2.
@@ 3 Results
For clarity and ease of comparison, all figures of observations in this section consistently present the 2023 July 13 event in the left-hand panels and the 2024 July 23 event in the right-hand panels. Figure 1 illustrates the two solar eruptions and the respective spacecraft locations analyzed in this study. On 2023 July 13, an active region was located on the far side of the Sun at N15E160 as viewed from Earth (upper left panel), while Solar Orbiter was positioned at Stonyhurst longitude. At approximately 06:46 UT, this region produced an intense X-ray flare, recorded by the Spectrometer and Telescope for Imaging X-rays (STIX;Krucker et al. 2020) on board Solar Orbiter, with a magnitude estimated to be X1 on the GOES scale. This eruption was accompanied by a halo CME observed by SOHO/LASCO (lower left panel), with the associated CME propagating at approximately km/s at a heliocentric distance of about , according to the DONKI database 111https://kauai.ccmc.gsfc.nasa.gov/DONKI/view/CMEAnalysis/25934/2. Similarly, for the 2024 July 23 event, Solar Orbiter was located at Stonyhurst longitude. An active region was present at S01W164, also on the far side of the Sun and toward Solar Orbiter (upper right panel). At around 23:56 UT on 2024 July 22, this region produced a powerful flare estimated to be X14 on the GOES scale. The corresponding halo CME (lower right panel) had an estimated speed of approximately km/s at a heliocentric distance of about 222https://kauai.ccmc.gsfc.nasa.gov/DONKI/view/CMEAnalysis/32125/2. We note that both events originated on the far side of the Sun, so the estimated CME speeds may be subject to significant uncertainty.
@@
Figure 2 provides an overview of the two SEP events examined in this study, hereafter referred to as event 1 (2023 July 13) and event 2 (2024 July 23), as observed by Solar Orbiter at heliocentric distances of and , respectively. We first focus here on event 1 as shown in the left-hand panels. The top bar indicates that the in-situ magnetic field direction is predominantly outward from the Sun during the interval of interest. Panel (a) presents the omni-directional proton dynamic spectra with a time resolution of , revealing an initial weaker particle injection followed by a much stronger injection spanning energies from a few MeV up to (for the dynamic spectra of the individual viewing direction, see Fig. 3). The black and magenta dashed lines indicate the fitted velocity dispersion tracks (more details are provided in Fig. 3). The first-order anisotropy , shown in panel (b), is a key diagnostic in this work. We derive the first-order anisotropy using directional intensity measurements from both EPT and HET, which both measure in four directions. The first-order anisotropy is defined as , where is the pitch angle dependent intensity in a given direction and is the average pitch angle cosine. Here we compute the anisotropy for each time and energy bin. The sign of is interpreted in the context of the magnetic field polarity: for a positive (outward) field, positive indicates anti-sunward particle flow, while negative corresponds to sunward flow. The interpretation is reversed for negative polarity. Given that the magnetic field is mostly outward in event 1, large positive anisotropy indicates that particles are streaming away from the Sun, whereas negative anisotropy denotes sunward streaming. A pronounced anisotropy (its magnitude close to 3) is observed at energies above , and its velocity dispersion suggests two separate injections within the first six hours following the event onset. Notably, the higher particle intensity corresponds to the sunward particle flow. Panel (c) displays the pitch angle distribution of protons, with the color scale representing normalized intensity, calculated as the ratio of the flux measured by each telescope to the mean flux across four telescopes. After the onset, the intensity is initially highest for pitch angles below (from the sunward telescope) and then shifts to above (anti-sunward telescope), indicating a stronger particle flow directed toward the Sun later on. Panel (d) shows the pitch angle distribution of protons. During the period of 9:00 to 10:00 UT, the protons show peak intensity with pitch angles , while the protons display maximum intensity with pitch angles . This comparison demonstrates that Solar Orbiter observed two populations from opposite directions simultaneously. Panel (e) presents the pitch angle distribution of electron measurements from SWA/EAS. A distinct bidirectional electron streaming is apparent from around 03:00 to 12:00 UT on 2023 July 13, as indicated by enhanced phase space density near pitch angles and and a clear loss cone around . This suggests that Solar Orbiter might be situated on a closed magnetic structure with both footpoints connected to the Sun. Based solely on the plasma and magnetic field measurements, it is not possible to visually confirm a classical magnetic flux rope between 03:00 and 12:00 UT due to the absence of clear rotations or significant plasma signatures from panels (f)-(i). To further characterize the possible magnetic structure, we apply the PyGS algorithm to identify possible SFRs during this interval. The resulting SFRs are indicated by the gray-shaded area, with a primary SFR detected between 03:21 and 08:45 UT. This flux rope is likely responsible for the observed bidirectional streaming particle in this event, although its start and end times do not fully coincide with the period of bidirectional electrons. As discussed in Section 1, the origin of SFRs may be either locally generated or associated with small CMEs. Given the relatively long duration of the SFR (approximately 5 hours) and the presence of bidirectional suprathermal electrons, we suggest that this structure likely corresponds to a small interplanetary CME. Panel (j) displays the normalized reduced magnetic helicity, with white contour lines enclosing regions where . These high magnetic-helicity regions are observed between around 05:00 and 08:00 UT, coinciding with the duration of the SFR identified by PyGS. We further focus on this interval and examine the normalized cross helicity and normalized residual energy , which suggest the presence of an Alfvénic SFR, as detailed in Appendix A.2.
@@
We attempt to trace the origin of the SFR back to the Sun. According to the DONKI database, multiple small CMEs erupted from the same active region as event 1 between 2023 July 8 and 2023 July 10. This scenario is also supported by the identification of three in-situ shocks on 2023 July 13 (Trotta et al., 2025), as indicated by the vertical dashed lines in panels (f)-(i). Assuming the speed of a small CME ranges from to as suggested by in-situ solar wind measurements, the propagation time from the Sun to Solar Orbiter is approximately to days This implies an eruption time between 2023 July 8, 10:00 UT and July 9, 00:00 UT, given an in-situ onset of SFR at around 2023 July 13, 03:00 UT. Within this window, only one candidate CME was recorded, launched on 2023 July 8, 20:36 UT, with a reported speed of and a half-width of 333https://kauai.ccmc.gsfc.nasa.gov/DONKI/view/CME/25858/1. The associated flare was located at N15W140 as viewed from Earth. Therefore, both the SFR and event 1 likely originated from the same active region. However, we do not investigate this further due to limited observational evidence.

### v2

@@ Abstract
Aims. We investigate the origin and transport of energetic particles in two rare bidirectional anisotropic SEP events observed by Solar Orbiter, with a particular emphasis on their association with magnetic flux ropes.
Methods. Energetic particles, solar wind plasma, magnetic field, and solar radio measurements were analysed. Via the velocity dispersion analysis, we determined release times and path lengths for distinct particle populations. Automated flux rope identification and magnetic helicity diagnostics were used to characterize magnetic flux ropes.
Results. Both events showed two clear velocity dispersion signatures with opposite particle anisotropies during their onset phase. The sunward streaming protons, characterized by a delayed release time, a harder spectral index, and higher intensities, may be due to coronal mass ejection-driven shock acceleration, while the promptly released anti-sunward streaming protons are likely linked to flare acceleration. Notably, in both cases, small-scale flux ropes were identified in situ during the time intervals corresponding to the bidirectional particle streaming. Path lengths derived for sunward and anti-sunward injections were substantially greater than nominal values of the Parker field lines, further evidence of the role of the flux rope in shaping particle trajectories.
Conclusions. These observations demonstrate that magnetic flux ropes can significantly affect magnetic connectivity to the source region and SEP propagation in the inner heliosphere, and that simultaneous velocity dispersion from two distinct particle sources can be used to place direct constraints on the topology of the flux rope. Our results highlight the value of combining particle anisotropy, release time, source spectra, and magnetic structure diagnostics to unravel SEP transport in complex transient magnetic structures, and also present new challenges for the current SEP transport model.
@@
Solar energetic particle (SEP) events are characterized by transient enhancements in the intensity of energetic charged particles, which are accelerated from the solar corona during eruptive phenomena such as flares and coronal mass ejections (CMEs) and subsequently propagate along magnetic field lines into interplanetary space (e.g. Desai & Giacalone, 2016). The particle anisotropy measured at the spacecraft serves as a direct indicator of particle transport processes. Specifically, a high anisotropy is typically associated with weak pitch angle scattering, meaning that the particle flux is strongly peaked along certain magnetic field directions, with most particles streaming either sunwards or anti-sunwards. In contrast, a low or near-zero anisotropy suggests strong pitch angle scattering or enhanced perpendicular diffusion, resulting in a more isotropic particle distribution (Dresing et al., 2014). Therefore, anisotropy is a crucial diagnostic for identifying atypical SEP transport processes, especially in complex interplanetary environments.
@@
While most SEP events exhibit outward streaming particles along open magnetic field lines, an increasing number of in situ observations have revealed rare cases of sunward-directed SEP events, where particles are detected propagating back towards the Sun with high intensity (e.g. Richardson & Cane, 1996; Larson et al., 1997; Rodriguez-Pacheco, 1999; Richardson et al., 2000; Malandraki et al., 2002; Rodr´ıguez-Pacheco et al., 2003; Leske et al., 2012; Tan et al., 2012; Gómez-Herrero et al., 2017; Li et al., 2020; Wei et al., 2024; Rodr´ıguez-Garc´ıa et al., 2025). These events typically involve special interplanetary conditions, such as an observer being embedded within a magnetic cloud (MC), the presence of a shock or compressed magnetic fields acting as a magnetic mirror, or remote acceleration sites away from the observer. Based on previous case studies, sunward SEP events can be generally categorized into four main types according to their underlying physical mechanisms. Firstly, the observer can reside within a closed magnetic flux rope (e.g. a MC) with both legs anchored at the Sun, enabling bidirectional (sunward and anti-sunward) particle flows via dual injections from two legs or reflection at the footpoint (Dresing et al., 2016; Gómez-Herrero et al., 2017; Rodr´ıguez-Garc´ıa et al., 2025). Moreover, the three-dimensional topology of the flux rope also allows for the possibility that particles are injected from the north or south directions (Wimmer-Schweingruber et al., 2023). Secondly, sunward SEPs can arise from magnetic mirroring, where SEPs move along open field lines and encounter a region of increased magnetic field strength, such as a shock or a CME sheath, which acts as a magnetic mirror and reflects particles towards the Sun (Malandraki et al., 2002; Klassen et al., 2012; Tan et al., 2012; Li et al., 2020). In contrast to cases where an observer inside a closed flux rope can detect SEPs from different sources at each footpoint, the mirror effect results in the observation of particles originating from the same source. Thirdly, Wei et al. (2024) propose that a remote strong accelerator (e.g. an interplanetary CME-driven shock) located beyond the observer can inject SEPs that then propagate sunwards along open magnetic field lines to the spacecraft. Finally, cross-field diffusion and field-line meandering may allow sunward SEPs to reach observers with poor magnetic connectivity to the source region (e.g. He, 2015). However, this mechanism alone is insufficient to explain sunward-propagating events observed at locations with good magnetic connectivity to the source.
@@
While previous studies have established several categories of sunward-propagating SEP events, here we investigate bidirectional SEP events that exhibit two distinct velocity dispersion signatures during their onset phases, one associated with anti-sunward propagation and the other with sunward propagation. Such events, although rare, can provide critical constraints on both the path lengths and release times of different particle sources and thereby offer unique insights into the large-scale configuration of the interplanetary magnetic field, as well as the particle acceleration and transport. In this work we present a detailed analysis of two bidirectional, anisotropic SEP events observed by Solar Orbiter, each displaying two distinct velocity dispersion signatures and both likely associated with small-scale flux ropes (SFRs). To our knowledge, such events have not been previously reported. SFRs typically have durations ranging from tens of minutes to several hours. The origin of SFRs remains a subject of debate, as they are believed to result from solar eruptive processes, possibly as manifestations of small CMEs (Feng et al., 2007), and from in situ formation in the heliosphere through magnetic reconnection (Cartwright & Moldwin, 2010). Notably, SFRs have much shorter durations and lack the typical plasma signatures of large-scale CMEs (e.g. Wimmer-Schweingruber et al., 2006; Hu et al., 2018). As a result, SFRs are more challenging to detect and interpret, yet they could play an important role in modulating energetic particle propagation in the solar wind.
@@
Through a combination of energetic particle measurements, velocity dispersion analysis (VDA), and magnetic field diagnostics, we investigated the origins, transport process, and associated magnetic structure of two bidirectional SEP events. Section 2 describes the data and methods employed. The observations and interpretation of the two events are presented in Sect. 3. Finally, Sect. 4 summarizes the main findings and discusses their implications for our understanding of SEP transport and the topology of flux ropes.
@@ 2 Data and methods
This study utilizes data collected by multiple instruments on board Solar Orbiter. Specifically, SEP measurements are provided by the Energetic Particle Detector (EPD) suite (Rodr´ıguez-Pacheco et al. 2020; Wimmer-Schweingruber et al. 2021), which comprises the Supra-Thermal Electron and Proton sensor (STEP), the Electron Proton Telescope (EPT), the High Energy Telescope (HET), and the Suprathermal Ion Spectrograph (SIS). These instruments cover ion energies from a few keV nucleon-1 up to above 100 MeV nucleon-1. In this work we mainly utilized ion measurements from EPT and HET; each offers four viewing directions, thereby enabling the analysis of particle anisotropy. While EPT does not explicitly differentiate between ion species, the measured fluxes are commonly interpreted as proton-dominated, since protons constitute the most abundant ion population in large SEP events. Magnetic field data are obtained from the Solar Orbiter Magnetometer (MAG; Horbury et al. 2020). Plasma parameters are measured with the Proton-Alpha Sensor (PAS), and the pitch angle distributions of suprathermal electrons are provided by the Electron Analyser System (EAS), both part of the Solar Wind Analyzer (SWA; Owen et al. 2020). Additionally, we examined radio data from the Radio and Plasma Waves (RPW; Maksimovic et al. 2020) instrument on board Solar Orbiter.
@@
A key focus of this work is to investigate the roles of small-scale magnetic flux ropes in the bidirectional SEP events. Identifying SFRs in the solar wind presents a significant challenge due to their short durations, weaker coherent magnetic signatures compared to MCs, and ambiguous boundaries, especially in the presence of solar wind fluctuations. To address this, we employed the automated detection algorithm implemented in the PyGS package (Hu et al., 2018; Zheng & Hu, 2018; Chen & Hu, 2022), which leverages the unique characteristics of the Grad–Shafranov (GS) equation (Sonnerup & Guo, 1996; Hu & Sonnerup, 2002). In the local flux rope frame, a spacecraft crossing an SFR typically observes a bipolar rotation in one of the magnetic field components, corresponding to nested isosurfaces of the magnetic flux function . For a true flux rope, the transverse pressure () is expected to be a single-valued function of , resulting in a characteristic double-folding pattern of between the inbound and outbound traversals. The PyGS algorithm systematically scans in situ data for intervals that exhibit this double-folding signature, enabling a robust and automated identification of SFRs.
@@
However, in highly dynamic or turbulent environments, conventional automated methods such as PyGS may fail to reliably identify flux rope structures due to strong magnetic field fluctuations. To address these limitations, we supplemented our analysis with calculations of the normalized reduced magnetic helicity, , which serves as a sensitive proxy for magnetic field twist. Following Zhao et al. (2020, 2021b), the scale- and time-dependent normalized reduced magnetic helicity, , is then evaluated as
@@
where the tilde represents wavelet-transformed quantities using the complex Morlet wavelet function (Torrence & Compo, 1998), refers to the imaginary part, represents the wavelet scale, and the asterisk denotes complex conjugation. Here, , , and correspond to the magnetic field components in the radial-tangential-normal (RTN) coordinate system. As demonstrated in previous studies (Zhao et al., 2020, 2021a, 2021b), flux ropes are typically characterized by high values of magnetic helicity, which result from the large-angle rotation of the magnetic field. In this work we used regions with as proxies for the possible presence of SFRs. Since strongly Alfvénic fluctuations also typically appear as intermittent enhancements of , to exclude their influence, we further examined the normalized cross helicity and the normalized residual energy , as described in Appendix A.2.
@@ 3 Results
For clarity and ease of comparison, all figures of observations in this section consistently present the July 13, 2023, event (event 1) in the left-hand panels and the July 23, 2024, event (event 2) in the right-hand panels. Figure 1 illustrates the two solar eruptions and the respective spacecraft locations analysed in this study. During event 1, an active region was located on the far side of the Sun at N15E160 as viewed from Earth (upper-left panel), while Solar Orbiter was positioned at Stonyhurst longitude. At approximately 06:46 UT, this region produced an intense X-ray flare, recorded by the Spectrometer and Telescope for Imaging X-rays (STIX;Krucker et al. 2020) on board Solar Orbiter, with a magnitude estimated to be X1 on the GOES scale. This eruption was accompanied by a halo CME observed by SOHO/LASCO (lower-left panel), with the associated CME propagating at approximately km/s at a heliocentric distance of about , according to the DONKI database 111https://kauai.ccmc.gsfc.nasa.gov/DONKI/view/CMEAnalysis/25934/2. Similarly, for event 2, Solar Orbiter was located at Stonyhurst longitude. An active region was present at S01W164, also on the far side of the Sun and towards Solar Orbiter (upper-right panel). At around 23:56 UT on July 22, 2024, this region produced a powerful flare estimated to be X14 on the GOES scale. The corresponding halo CME (lower-right panel) had an estimated speed of approximately km/s at a heliocentric distance of about 222https://kauai.ccmc.gsfc.nasa.gov/DONKI/view/CMEAnalysis/32125/2. We note that both events originated on the far side of the Sun, so the estimated CME speeds may be subject to significant uncertainty.
@@
Figure 2 provides an overview of the two SEP events examined in this study, event 1 and event 2, as observed by Solar Orbiter at heliocentric distances of and , respectively. We first focus on event 1, as shown in the left-hand panels. The top bar indicates that the in situ magnetic field direction is predominantly outwards from the Sun during the interval of interest. Panel (a) presents the omnidirectional proton dynamic spectra with a time resolution of , revealing an initial weaker particle injection followed by a much stronger injection spanning energies from a few MeV up to (for the dynamic spectra of the individual viewing direction, see Fig. 3). The black and magenta dashed lines indicate the fitted velocity dispersion tracks (more details are provided in Fig. 3). The first-order anisotropy , shown in panel (b), is a key diagnostic in this work. We derived the first-order anisotropy using directional intensity measurements from both EPT and HET, which both measure in four directions. The first-order anisotropy is defined as , where is the pitch-angle-dependent intensity in a given direction and is the average pitch angle cosine. Here we computed the anisotropy for each time and energy bin. The sign of is interpreted in the context of the magnetic field polarity: for a positive (outward) field, positive indicates anti-sunward particle flow, while negative corresponds to sunward flow. The interpretation is reversed for negative polarity. Given that the magnetic field is mostly outwards in event 1, large positive anisotropy indicates that particles are streaming away from the Sun, whereas negative anisotropy denotes sunward streaming. A pronounced anisotropy (its magnitude close to 3) is observed at energies above , and its velocity dispersion suggests two separate injections within the first six hours following the event onset. Notably, the higher particle intensity corresponds to the sunward particle flow. Panel (c) displays the pitch angle distribution of protons, with the colour scale representing normalized intensity, calculated as the ratio of the flux measured by each telescope to the mean flux across four telescopes. After the onset, the intensity is initially highest for pitch angles below (from the sunward telescope) and then shifts to above (anti-sunward telescope), indicating a stronger particle flow directed towards the Sun later on. Panel (d) shows the pitch angle distribution of protons. During the period of 9:00 to 10:00 UT, the protons show peak intensity with pitch angles , while the protons display maximum intensity with pitch angles . This comparison demonstrates that Solar Orbiter observed two populations from opposite directions simultaneously. Panel (e) presents the pitch angle distribution of electron measurements from SWA/EAS. A distinct bidirectional electron streaming is apparent from around 03:00 to 12:00 UT on July 13, 2023, as indicated by enhanced phase space density near pitch angles and . This suggests that Solar Orbiter might be situated on a closed magnetic structure with both footpoints connected to the Sun. Based solely on the plasma and magnetic field measurements, it is not possible to visually confirm a classical magnetic flux rope between 03:00 and 12:00 UT due to the absence of clear rotations or significant plasma signatures from panels (f)-(i). To further characterize the possible magnetic structure, we applied the PyGS algorithm to identify possible SFRs during this interval. The resulting SFRs are indicated by the grey-shaded area, with a primary SFR detected between 03:21 and 08:45 UT. This flux rope is likely responsible for the observed bidirectional streaming particle in this event, although its start and end times do not fully coincide with the period of bidirectional electrons. As discussed in Sect. 1, the origin of SFRs can be either locally generated or associated with small CMEs. Given the relatively long duration of the SFR (approximately 5 hours) and the presence of bidirectional suprathermal electrons, we suggest that this structure likely corresponds to a small interplanetary CME. Panel (j) displays the normalized reduced magnetic helicity, with white contour lines enclosing regions where . These high magnetic-helicity regions are observed between around 05:00 and 08:00 UT, coinciding with the duration of the SFR identified by PyGS. We further focused on this interval and examined the normalized cross helicity () and normalized residual energy (), which suggest the presence of an Alfvénic SFR, as detailed in Appendix A.2.
@@
We attempted to trace the origin of the SFR back to the Sun. According to the DONKI database, multiple small CMEs erupted from the same active region as event 1 between July 8, 2023, and July 10, 2023. This scenario is also supported by the identification of three in situ shocks on July 13, 2023 (Trotta et al., 2025), as indicated by the vertical dashed lines in panels (f)-(i). Assuming the speed of a small CME ranges from to as suggested by in situ solar wind measurements, the propagation time from the Sun to Solar Orbiter is approximately to days This implies an eruption time between July 8, 2023, at 10:00 UT and July 9, at 00:00 UT, given an in situ onset of SFR at around 03:00 UT on July 13, 2023. Within this window, only one candidate CME was recorded, launched on July 8, 2023, at 20:36 UT, with a reported speed of and a half-width of 333https://kauai.ccmc.gsfc.nasa.gov/DONKI/view/CME/25858/1. The associated flare was located at N15W140 as viewed from Earth. Therefore, the SFR and event 1 likely originated from the same active region. However, we did not investigate this further due to limited observational evidence.